# Distribution Estimation with Side Information

### v1

@@ III-B Model 2
With the side information that $S_{l}$ is partitioned into $S_{l}^{A}$ and $S_{l}^{B}$, the two-level estimator we provide treats the problem as two separate distribution estimation problems: one on the symbols in $A$ and the other on symbols in $B$. We perform Good-Turing estimation on the elements in $A$ using $N_{A}$ samples and Good-Turing estimation on the elements in $B$ using $N_{B}$ samples. Thus, our two-level estimate is
$\hat{\pi}^{(l,A)}=\frac{l+1}{n-l}\frac{\phi_{l+1}^{A}}{\phi_{l}^{A}},\qquad\hat{\pi}^{(l,B)}=\frac{l+1}{n-l}\frac{\phi_{l+1}^{B}}{\phi_{l}^{B}}.$ (21)
If the probabilities in $A$ and $B$ are sufficiently apart, the two-level estimator is expected to perform well. However, if the true distribution is close to the uniform distribution, then the two-level estimate is expected to perform worse since we are using fewer samples (i.e. $N_{A}$ or $N_{B}$ instead of $n$) to estimate probabilities in each partition. This is a trade-off inherent to the problem.
For the two-level estimator in (21), observe that
$\phi_{l}\hat{\pi}^{(l)}=\phi_{l}^{A}\hat{\pi}^{(l,A)}+\phi_{l}^{B}\hat{\pi}^{(l,B)}.$ (22)
Now, we characterize how much better the two-level estimator performs as compared to the one-level estimator.
We first note the following algebraic identity.
@@
$=$ $\frac{\phi_{l}^{A}\phi_{l}^{B}}{\phi_{l}}\Bigg[(\pi^{\textsf{(or, A)}}-\pi^{\textsf{(or, B)}})^{2}$
$-\left((\hat{\pi}^{(l,A)}-\pi^{\textsf{(or, A)}})-(\hat{\pi}^{(l,B)}-\pi^{\textsf{(or, B)}})\right)^{2}\Bigg].$
The partition-wise Good-Turing estimators ensure that the value of the $\left(\hat{\pi}^{(l,A)}-\pi^{\textsf{(or, A)}}-\hat{\pi}^{(l,B)}+\pi^{\textsf{(or, B)}}\right)^{2}$ term would be small in expectation. In particular, we note the following theorem.
@@
For $l\leq\frac{n}{4}\min(\pi_{A},\pi_{B})$, we have
$\mathbb{E}\left[\frac{\phi_{l}^{A}\phi_{l}^{B}}{\phi_{l}}\left(\hat{\pi}^{(l,A)}-\pi^{\textsf{(or, A)}}-\hat{\pi}^{(l,B)}+\pi^{\textsf{(or, B)}}\right)^{2}\right]\\ \leq\frac{2c_{l}}{n}\left(\frac{1}{\pi_{A}}+\frac{1}{\pi_{B}}\right).$ (25)
Thus, for sufficiently large $n$, the two-level estimator has a gain of
$\mathbb{E}\left[\frac{\phi_{l}^{A}\phi_{l}^{B}}{\phi_{l}}(\pi^{\textsf{(or, A)}}-\pi^{\textsf{(or, B)}})^{2}\right]$ (26)
over the one-level estimator for species in $S_{l}$.
Remarks: (1) While we assume initially that $A$ and $B$ are both known, it is sufficient if we know the sets $S_{l}^{A}$, $S_{l}^{B}$, $S_{l+1}^{A}$ and $S_{l+1}^{B}$. (2) Lower bounds for the partial order side information is complicated by the occurrence of $\phi_{l}$ in (26) and will be taken up as future work.
@@ VI-B Proof of Theorem 15
    $\max_{v,v^{\prime}:d_{H}(v,v^{\prime})=1}d_{KL}(\theta_{v},\theta_{v^{\prime}}).$
@@ VI-D Proof of Theorem 17
We pick $\tau=\min(\frac{\Delta}{\sqrt{12}},\frac{1}{2\sqrt{n}})$ to get the desired result.
@@
Observe that
$Q_{l}(\pi,\hat{\pi}^{(l)})$ $=$ $\sum_{i\in S_{l}}(\pi_{i}-\hat{\pi}^{(l)})^{2}$
$=$ $\sum_{i\in S_{l}^{A}}(\pi_{i}-\hat{\pi}^{(l)})^{2}+\sum_{i\in S_{l}^{B}}(\pi_{i}-\hat{\pi}^{(l)})^{2}$
$\overset{(a)}{=}$ $\sum_{i\in S_{l}^{A}}(\pi_{i}-\pi^{\textsf{(or, A)}})^{2}+\phi_{l}^{A}(\pi^{\textsf{(or, A)}}-\hat{\pi}^{(l)})^{2}$
$\quad+\sum_{i\in S_{l}^{B}}(\pi_{i}-\pi^{\textsf{(or, B)}})^{2}+\phi_{l}^{B}(\pi^{\textsf{(or, B)}}-\hat{\pi}^{(l)})^{2},$
where $(a)$ follows by adding and subtracting $\pi^{\textsf{(or, A)}}$ and $\pi^{\textsf{(or, B)}}$ to the respective terms. Further using (19), we get
@@
$=$ $\phi_{l}^{A}(\pi^{\textsf{(or, A)}}-\hat{\pi}^{(l)})^{2}+\phi_{l}^{B}(\pi^{\textsf{(or, B)}}-\hat{\pi}^{(l)})^{2}$
@@
Now note that
$\phi_{l}^{A}((\pi^{\textsf{(or, A)}}-\hat{\pi}^{(l)})^{2}-(\pi^{\textsf{(or, A)}}-\hat{\pi}^{(l,A)})^{2})\\ =\phi_{l}^{A}((\hat{\pi}^{(l)})^{2}-(\hat{\pi}^{(l,A)})^{2}+2\pi^{\textsf{(or, A)}}(\hat{\pi}^{(l,A)}-\hat{\pi}^{(l)}))$ (48)
and that a similar identity exists for the set $B$ as well. Using this,
@@
$=$ $\phi_{l}^{A}((\hat{\pi}^{(l)})^{2}-(\hat{\pi}^{(l,A)})^{2}+2\pi^{\textsf{(or, A)}}(\hat{\pi}^{(l,A)}-\hat{\pi}^{(l)}))$
$+\phi_{l}^{B}((\hat{\pi}^{(l)})^{2}-(\hat{\pi}^{(l,B)})^{2}+2\pi^{\textsf{(or, B)}}(\hat{\pi}^{(l,B)}-\hat{\pi}^{(l)}))$
@@
$+\phi_{l}^{A}(-(\hat{\pi}^{(l,A)})^{2}+2\pi^{\textsf{(or, A)}}\hat{\pi}^{(l,A)})$
$+\phi_{l}^{B}(-(\hat{\pi}^{(l,B)})^{2}+2\pi^{\textsf{(or, B)}}\hat{\pi}^{(l,B)}).$
Next, we get
[…]
$\overset{(a)}{=}$ $\frac{(\phi_{l}^{A}\hat{\pi}^{(l,A)})^{2}+(\phi_{l}^{B}\hat{\pi}^{(l,B)})^{2}+2\phi_{l}^{A}\phi_{l}^{B}\hat{\pi}^{(l,A)}\hat{\pi}^{(l,B)}}{\phi_{l}}$
$-2\frac{(\phi_{l}^{A}\hat{\pi}^{(l,A)}+\phi_{l}^{B}\hat{\pi}^{(l,B)})}{\phi_{l}}(\pi^{\textsf{(or, A)}}\phi_{l}^{A}+\pi^{\textsf{(or, B)}}\phi_{l}^{B})$
$+\phi_{l}^{A}(-(\hat{\pi}^{(l,A)})^{2}+2\pi^{\textsf{(or, A)}}\hat{\pi}^{(l,A)})$
$+\phi_{l}^{B}(-(\hat{\pi}^{(l,B)})^{2}+2\pi^{\textsf{(or, B)}}\hat{\pi}^{(l,B)})$
$\overset{(b)}{=}$ $\frac{-\phi_{l}^{A}\phi_{l}^{B}(\hat{\pi}^{(l,A)})^{2}}{\phi_{l}}+\frac{-\phi_{l}^{A}\phi_{l}^{B}(\hat{\pi}^{(l,B)})^{2}}{\phi_{l}}$
$+\frac{2\phi_{l}^{A}\phi_{l}^{B}\hat{\pi}^{(l,A)}\hat{\pi}^{(l,B)}}{\phi_{l}}$
$+2\frac{\phi_{l}^{A}\hat{\pi}^{(l,A)}\pi^{\textsf{(or, A)}}\phi_{l}^{B}+\phi_{l}^{A}\hat{\pi}^{(l,B)}\pi^{\textsf{(or, B)}}\phi_{l}^{B}}{\phi_{l}}$
$-2\frac{\phi_{l}^{B}\hat{\pi}^{(l,B)}\pi^{\textsf{(or, A)}}\phi_{l}^{A}+\phi_{l}^{A}\hat{\pi}^{(l,A)}\pi^{\textsf{(or, B)}}\phi_{l}^{B}}{\phi_{l}}$
$\overset{(c)}{=}$ $\frac{\phi_{l}^{A}\phi_{l}^{B}}{\phi_{l}}\bigg[-(\hat{\pi}^{(l,A)})^{2}-(\hat{\pi}^{(l,B)})^{2}$
$+2\hat{\pi}^{(l,A)}\hat{\pi}^{(l,B)}+2\hat{\pi}^{(l,A)}\pi^{\textsf{(or, A)}}$
$+2\hat{\pi}^{(l,B)}\pi^{\textsf{(or, B)}}-2\hat{\pi}^{(l,B)}\pi^{\textsf{(or, A)}}-2\hat{\pi}^{(l,A)}\pi^{\textsf{(or, B)}}\bigg],$
where $(a)$ uses $\hat{\pi}^{(l)}=\dfrac{(\phi_{l}^{A}\hat{\pi}^{(l,A)}+\phi_{l}^{B}\hat{\pi}^{(l,B)})}{\phi_{l}}$, $(b)$ gathers all terms and uses $\phi_{l}=\phi_{l}^{A}+\phi_{l}^{B}$ repeatedly, and $(c)$ takes the $\phi_{l}^{A}\phi_{l}^{B}$ term as a common factor.
It can be seen that this is equal to the term given in the lemma statement.
VI-F Proof of Theorem 25
First, observe that
$\mathbb{E}\left[\frac{\phi_{l}^{A}\phi_{l}^{B}}{\phi_{l}}\left(\hat{\pi}^{(l,A)}-\pi^{\textsf{(or, A)}}-\hat{\pi}^{(l,B)}+\pi^{\textsf{(or, B)}}\right)^{2}\right]\\ \leq\mathbb{E}\left[\frac{\phi_{l}^{A}\phi_{l}^{B}}{\phi_{l}}((\hat{\pi}^{(l,A)}-\pi^{\textsf{(or, A)}})^{2}+(\hat{\pi}^{(l,B)}-\pi^{\textsf{(or, B)}})^{2})\right]$ (49)
Next, out of the $n$ samples, let $N_{A}$ and $N_{B}$ be the number of samples from the sets $A$ and $B$, respectively. Clearly, $N_{A}+N_{B}=n$ and $N_{A}\sim\textsf{Bin}(n,\pi_{A})$ and $N_{B}\sim\textsf{Bin}(n,\pi_{B})$. Now, note that
[…]
$\overset{(a)}{\leq}$ $\mathbb{E}\left[(\phi_{l}^{A})^{2}\cdot(\hat{\pi}^{(l,A)}-\pi^{\textsf{(or, A)}})^{2}\right],$
$=$ $\mathbb{E}[(M_{l}^{A}-\hat{M}_{l}^{A})^{2}]$
where $(a)$ uses $\frac{\phi_{l}^{A}\phi_{l}^{B}}{\phi_{l}}\leq\phi_{l}^{A}\leq(\phi_{l}^{A})^{2}$. Next, note from Theorem II.2 that
$\mathbb{E}[(M_{l}^{A}-\hat{M}_{l}^{A})^{2}|N_{A}=n_{A}]\leq\left(\frac{c_{l}}{n_{A}}\right).$
Using $(M_{l}^{A}-\hat{M}_{l}^{A})^{2}\leq 1$, we get a looser bound:
$\mathbb{E}[(M_{l}^{A}-\hat{M}_{l}^{A})^{2}|N_{A}=n_{A}]\leq\left(\frac{c_{l}}{n_{A}+1}\right).$
From this, note that
$\mathbb{E}[(M_{l}^{A}-\hat{M}_{l}^{A})^{2}]$ $\overset{(a)}{\leq}$ $\mathbb{E}\left[\frac{2c_{l}}{N_{A}+1}\right]$
$\overset{(b)}{=}$ $2c_{l}\frac{(1-(1-\pi_{A})^{n+1})}{(n+1)\pi_{A}}$
$\overset{(c)}{\leq}$ $2\frac{c_{l}}{n\pi_{A}},$
where $(a)$ uses the theorem of total probability, $(b)$ uses standard expected value calculations, and $(c)$ uses $(1-\pi_{A})^{n+1}\geq 0$ and $\frac{1}{n+1}\leq\frac{1}{n}$.
Applying a similar procedure for the set $B$ gives the theorem directly.

### v2

@@ III-B Model 2
The quantity of interest to us is $Q_{l}(\pi,\hat{\pi}^{(l)})-Q_{l}(\pi,\hat{\pi}^{(l,A)},\hat{\pi}^{(l,B)})$, which in some sense is a regret quantity measuring the improvement in performance of the two-level estimator in comparison to the one-level estimator. We note the following algebraic lemma, which crucially uses (20).
@@
$=$ $\frac{\phi_{l}^{A}\phi_{l}^{B}}{\phi_{l}}(\pi^{\textsf{(or, A)}}-\pi^{\textsf{(or, B)}})^{2}+\phi_{l}(\pi^{(\textsf{or})}-\hat{\pi}^{(l)})^{2}$
$-\phi_{l}^{A}(\pi^{\textsf{(or, A)}}-\hat{\pi}^{(l,A)})^{2}-\phi_{l}^{B}(\pi^{\textsf{(or, B)}}-\hat{\pi}^{(l,B)})^{2}.$
We would want the $\phi_{l}^{A}(\pi^{\textsf{(or, A)}}-\hat{\pi}^{(l,A)})^{2}$ and $\phi_{l}^{B}(\pi^{\textsf{(or, B)}}-\hat{\pi}^{(l,B)})^{2}$ terms to be ‘small’. With the side information that $S_{l}$ is partitioned into $S_{l}^{A}$ and $S_{l}^{B}$, the two-level estimator we provide treats the problem as two separate distribution estimation problems: one on the symbols in $A$ and the other on symbols in $B$. We perform Good-Turing estimation on the elements in $A$ using $N_{A}$ samples and Good-Turing estimation on the elements in $B$ using $N_{B}$ samples. Thus, our two-level estimate is
$\hat{\pi}^{(l,A)}$ $=$ $\frac{l+1}{\max(1,N_{A}-l)}\frac{\phi_{l+1}^{A}}{\phi_{l}^{A}}$ (22)
$\hat{\pi}^{(l,B)}$ $=$ $\frac{l+1}{\max(1,N_{B}-l)}\frac{\phi_{l+1}^{B}}{\phi_{l}^{B}}.$ (23)
We note the following theorem on Good-Turing estimation of the two-level error terms.
@@
For $l\leq\frac{n}{2}\min(\pi_{A},\pi_{B})$, we have
$\mathbb{E}\left[\phi_{l}^{A}(\pi^{\textsf{(or, A)}}-\hat{\pi}^{(l,A)})^{2}+\phi_{l}^{B}(\pi^{\textsf{(or, B)}}-\hat{\pi}^{(l,B)})^{2}\right]\\ \leq\frac{c^{\prime}_{l}}{n}\left(\frac{1}{\pi_{A}}+\frac{1}{\pi_{B}}\right),$ (24)
where $c^{\prime}_{l}=8l+2c_{l}$ is a constant which depends only on $l$.
If the probabilities in $A$ and $B$ are sufficiently apart and $n$ is sufficiently large, the two-level estimator is expected to perform well. However, if the true distribution is close to the uniform distribution, then the two-level estimate is expected to perform worse since we are using fewer samples (i.e. $N_{A}$ or $N_{B}$ instead of $n$) to estimate probabilities in each partition. This is a trade-off inherent to the problem.
Note from using $(\pi^{(\textsf{or})}-\hat{\pi}^{(l)})^{2}\geq 0$ in Lemma III.5 and Theorem III.6 that
[…]
$\geq$ $\mathbb{E}\left[\frac{\phi_{l}^{A}\phi_{l}^{B}}{\phi_{l}}(\pi^{\textsf{(or, A)}}-\pi^{\textsf{(or, B)}})^{2}\right]-\frac{c^{\prime}_{l}}{n}\left(\frac{1}{\pi_{A}}+\frac{1}{\pi_{B}}\right).$
From the above, an important component of the expected gain of the two-level estimator over the one-level estimator for species in $S_{l}$ is seen to be
$\mathbb{E}\left[\frac{\phi_{l}^{A}\phi_{l}^{B}}{\phi_{l}}(\pi^{\textsf{(or, A)}}-\pi^{\textsf{(or, B)}})^{2}\right].$ (26)
Remarks: (1) While we assume initially that $A$ and $B$ are both known, it is sufficient if we know the quantities $S_{l}^{A}$, $S_{l}^{B}$, $S_{l+1}^{A}$, $S_{l+1}^{B}$, $N_{A}$, and $N_{B}$. (2) Lower bounds for the partial order side information is complicated by the occurrence of $\phi_{l}$ in (26) and will be taken up as future work.
@@ VI-B Proof of Theorem 15
    $\max_{v,v^{\prime}:d_{H}(v,v^{\prime})=1}d_{KL}(\theta_{v},\theta_{v^{\prime}}).$ (39)
@@ VI-D Proof of Theorem 17
We pick $\tau=\min\left(\frac{\Delta}{\sqrt{12}},\frac{1}{2\sqrt{n}}\right)$ to get the desired result.
@@
First, note from (II-B2) and (19) that
@@
$=$ $\sum_{i\in S_{l}}(\pi_{i}-\pi^{(\textsf{or})})^{2}+\phi_{l}(\pi^{(\textsf{or})}-\hat{\pi}^{(l)})^{2}$
$-\sum_{i\in S_{l}^{A}}(\pi_{i}-\pi^{\textsf{(or, A)}})^{2}-\sum_{i\in S_{l}^{B}}(\pi_{i}-\pi^{\textsf{(or, B)}})^{2}$
$-\phi_{l}^{A}(\pi^{\textsf{(or, A)}}-\hat{\pi}^{(l,A)})^{2}-\phi_{l}^{B}(\pi^{\textsf{(or, B)}}-\hat{\pi}^{(l,B)})^{2}.$
We will first bound the oracle error terms. Note from (II-B2) that
$\sum_{i\in S_{l}^{A}}\left((\pi_{i}-\pi^{(\textsf{or})})^{2}-(\pi_{i}-\pi^{\textsf{(or, A)}})^{2}\right)=\phi_{l}^{A}(\pi^{\textsf{(or, A)}}-\pi^{(\textsf{or})})^{2}$
and that a similar decomposition follows for the elements in $S_{l}^{B}$ also. From this, we get
@@
$=$ $\phi_{l}^{A}(\pi^{\textsf{(or, A)}}-\pi^{(\textsf{or})})^{2}+\phi_{l}^{B}(\pi^{\textsf{(or, B)}}-\pi^{(\textsf{or})})^{2}.$
Next, consider $\phi_{l}^{A}(\pi^{\textsf{(or, A)}}-\pi^{(\textsf{or})})^{2}$. Note from (20) and $\phi_{l}=\phi_{l}^{A}+\phi_{l}^{B}$ that $\pi^{\textsf{(or, A)}}-\pi^{(\textsf{or})}=\pi^{\textsf{(or, A)}}-\frac{\phi_{l}^{A}\pi^{\textsf{(or, A)}}+\phi_{l}^{B}\pi^{\textsf{(or, B)}}}{\phi_{l}}=\frac{\phi_{l}^{B}}{\phi_{l}}(\pi^{\textsf{(or, A)}}-\pi^{\textsf{(or, B)}})$. From this, we get
$\phi_{l}^{A}(\pi^{\textsf{(or, A)}}-\pi^{(\textsf{or})})^{2}$ $=$ $\frac{(\phi_{l}^{B})^{2}\phi_{l}^{A}}{\phi_{l}^{2}}(\pi^{\textsf{(or, A)}}-\pi^{\textsf{(or, B)}})^{2}.$
Using this expression in (49) we get
@@
$=$ $\frac{(\phi_{l}^{B})^{2}\phi_{l}^{A}}{\phi_{l}^{2}}(\pi^{\textsf{(or, A)}}-\pi^{\textsf{(or, B)}})^{2}+\frac{(\phi_{l}^{A})^{2}\phi_{l}^{B}}{\phi_{l}^{2}}(\pi^{\textsf{(or, A)}}-\pi^{\textsf{(or, B)}})^{2}$
$=$ $\frac{\phi_{l}^{A}\phi_{l}^{B}}{\phi_{l}}(\pi^{\textsf{(or, A)}}-\pi^{\textsf{(or, B)}})^{2}$
Substituting this in (48) gives the lemma directly.
VI-F Proof of Theorem III.6
First, observe that
$\mathbb{E}[\phi_{l}^{A}(\pi^{\textsf{(or, A)}}-\hat{\pi}^{(l,A)})^{2}]\\ \leq\mathbb{E}[(\phi_{l}^{A})^{2}(\pi^{\textsf{(or, A)}}-\hat{\pi}^{(l,A)})^{2}]\leq\mathbb{E}[(M_{l}^{A}-\hat{M}_{l}^{A})^{2}],$
where $\hat{M}_{l}^{A}=\frac{l+1}{\max(1,N_{A}-l)}\phi_{l+1}^{A}=\phi_{l}^{A}\hat{\pi}^{(l,A)}$.
Out of the $n$ samples, note that $N_{A}$ and $N_{B}$ are the number of samples from the sets $A$ and $B$ respectively. Clearly, $N_{A}+N_{B}=n$ and $N_{A}\sim\textsf{Bin}(n,\pi_{A})$ and $N_{B}\sim\textsf{Bin}(n,\pi_{B})$. The remainder of the proof bounds $\mathbb{E}[(M_{l}^{A}-\hat{M}_{l}^{A})^{2}]$.
First, note from Theorem II.2 and $(l+1)\phi_{l+1}^{A}\leq N_{A}$ that
$\mathbb{E}[(M_{l}^{A}-\hat{M}_{l}^{A})^{2}|N_{A}=n_{A}]\leq\begin{cases}\displaystyle\frac{c_{l}}{n_{A}},&n_{A}\geq l+1\\ n_{A},&\text{otherwise}.\end{cases}$
A slightly looser bound yields
$\mathbb{E}[(M_{l}^{A}-\hat{M}_{l}^{A})^{2}|N_{A}=n_{A}]\leq\begin{cases}\displaystyle\frac{2c_{l}}{n_{A}+1},&n_{A}\geq l+1\\ n_{A},&\text{otherwise}.\end{cases}$ (51)
Now, note that
@@
$\overset{(a)}{=}$ $\sum_{n_{A}=0}^{n}\mathbb{E}[(M_{l}^{A}-\hat{M}_{l}^{A})^{2}|N_{A}=n_{A}]\,\mathbb{P}(N_{A}=n_{A})$
$\overset{(b)}{\leq}$ $l\,\mathbb{P}(N_{A}\leq l)+\mathbb{E}\left[\frac{2c_{l}}{N_{A}+1}\right]$
$\overset{(c)}{\leq}$ $l\,e^{-\frac{n\pi_{A}}{8}}+\mathbb{E}\left[\frac{2c_{l}}{N_{A}+1}\right]$
$\overset{(d)}{\leq}$ $\frac{8l}{n\pi_{A}}+2c_{l}\frac{(1-(1-\pi_{A})^{n+1})}{(n+1)\pi_{A}}$
$\overset{(e)}{\leq}$ $\frac{8l}{n\pi_{A}}+2c_{l}\frac{1}{n\pi_{A}}\overset{(f)}{\leq}\frac{c^{\prime}_{l}}{n\pi_{A}}$
where $(a)$ follows from the theorem of total probability, $(b)$ uses (51), $(c)$ follows from $l\leq\frac{n\pi_{A}}{2}$ in the Chernoff bound, $(d)$ uses $e^{-x}\leq\frac{1}{x}$ and standard expected value calculations, $(e)$ uses $(1-\pi_{A})^{n+1}\geq 0$ and $\frac{1}{n+1}\leq\frac{1}{n}$, and $(f)$ defines $c^{\prime}_{l}=8l+2c_{l}$ to be some constant.
Applying a similar procedure for the set $B$ gives the desired theorem.